# Tight contact structures without symplectic fillings are everywhere

### 27 Feb 2024 v4

@@ 5.2. Abundance of exotic structures on spheres
When $n\geq 5$, there exists a fillable contact structure $(\mathbb{S}^{2n+1},\xi^{\prime})$ homotopic to the standard almost contact structure, which is $1$-ADC, and such that $0<\dim SH^{2n-3}_{+}(\mathbb{S}^{2n+1},\xi^{\prime})<\infty$.
Proof.
@@
We consider the Brieskorn sphere $Y:=\Sigma(n,\ldots,n,n+1,p)$, where $p$ is a prime number with $p\gg n$. By [KvK16, (14)] or [Zho22c, (5.12)], the minimal Conley-Zehnder index of a small perturbation of the standard contact form on $Y$ is $4-n$ (obtained with $|I|=n+2$, $|I_{T}|=n$, $N=1$, $T=n$ in [Zho22c, (5.12)]). The SFT degrees are then at least $2$. The minimal Conley-Zehnder indices of other Morse-Bott families of Reeb orbits of the natural contact form (after suitable perturbation) is $6-n$ (given by $N=2$, $T=n$ in [Zho22c, (5.12)]) as $p\gg n$, and so they have SFT degree $4$. As a consequence, we have that $Y$ is $1$-ADC. Moreover, by the Morse-Bott spectral sequence and the index gap above, we have $\dim SH_{+}^{(n+1)-(4-n)}(Y)=\dim SH_{+}^{2n-3}(Y)=1$, solely contributed to by the family with the minimal Conley-Zehnder index. Now, by [KvK16, Proposition 3.6], $Y$ is homeomorphic to a sphere. Then there exists $M$, such that the $M$-th iterated self connected sum #$\#^{M}Y$ is diffeomorphic to the standard sphere. Since $\dim Y\geq 11$, the contact sum preserves the $1$-ADC property and $\dim SH_{+}^{2n-3}(\#^{M}Y)=\dim\bigoplus^{M}SH^{2n-3}_{+}(Y)=M$ by [CO18, Theorem 7.1, Proposition 9.19]. To obtain the standard almost contact structure, we choose $Y^{\prime}$ to be the flexibly fillable sphere with the opposite homotopy class  [^5] To be precise, here “opposite” is to be interpreted as the inverse w.r.t. the natural group structure on the space of almost contact structures on the sphere given by the connected sum operation. of almost contact structures as that of #$\#^{M}Y$, as in [Laz20a, Corollary 1.6]. Then $Y^{\prime}\#^{M}Y$ is a standard sphere with the standard almost contact structure, and it is $1$-ADC. Moreover, $SH^{2n-3}_{+}(Y^{\prime})=0$ when $n\geq 5$. Hence we may take $Y^{\prime}\#^{M}Y$ as the desired contact sphere. ∎
@@
One first realises the given almost contact structure as the contact boundary $(M,\xi_{flex})$ of a flexible Weinstein domain using Eliashberg’s h-principle [CE12]. For $i\in\mathbb{N}$, we consider the iterated connected sums $Y_{i}:=(\mathbb{S}^{2n+1},\xi_{ex})\#(M,\xi_{flex})\#^{i}(\mathbb{S}^{2n+1},\xi^{\prime})$ with the contact sphere $(\mathbb{S}^{2n+1},\xi_{ex})$ from Theorem 5.1 and $(\mathbb{S}^{2n+1},\xi^{\prime})$ from Proposition 5.5. Notice that $M$ is $1$-ADC, as it is flexibly fillable [Laz20a], and that $SH^{*}_{+}(M)$ is supported in degree smaller than $n+1$, in view of the vanishing of symplectic homology for flexible Weinstein domains. Since $SH_{+}^{*}(\mathbb{S}^{2n+1},\xi_{ex})$ is supported in degrees at most $n+1$ by Lemma 5.4, then if $12n-3>n+$ 1, i.e. $n\geq 5$, we have $\dim SH^{2n-3}_{+}(Y_{i})=N\cdot i$, where $N=\dim SH^{2n-3}_{+}(\mathbb{S}^{2n+1},\xi^{\prime})$, by [CO18, Theorem 7.1, Proposition 9.19]. As a consequence, $Y_{i}$ are pairwise distinguished by $SH^{*}_{+}$. They are tight and have no strong filling by the same arguments as in the proof of Theorem A. ∎
@@
When $n=3$, $(\mathbb{S}^{2n+1},\xi_{ex})$ is not $1$-ADC, which costs us the well-definedness of the positive symplectic cohomology. When $n=4$, it is not easy to find a degree such that positive symplectic cohomology is different for varying summands in the connected sum. In those cases, we can use positive symplectic cohomology for augmentations in [Zho23, §3.2.3] [^6] The positive symplectic cohomology of $1$-ADC manifolds is the special case for the trivial augmentation., which was motivated by the work of Bourgeois-Oancea [BO09]. Then we can obtain a “contact invariant” by numerating through all possible augmentations. However, the current status of the contact homology [Par19, BH23] can not deduce that such contact invariant is independent of various choices. Assume the invariance of such theory, one can upgrade Theorem D to cover $n=3,4$.
@@ 6. Liouville fillable contact spheres without Weinstein fillings
We now explain how to use surgery techniques to deduce the existence of Liouville fillable contact structures that are not Weinstein fillable on spheres. This requires a refinement of the surgery arguments used above and we begin with the following:
@@
Let $(M^{2n+1},\xi)$ be a contact manifold of dimension at least $5$ whose underlying almost contact structure is stably trivial (as an almost contact structure). Then there is an almost complex bordism $W$ from $(M,\xi)$ to a homotopy sphere $\Sigma$ (with some almost contact structure).
@@
This follows immediately from results in [BCS14], building on work of Kreck [Kre99]. For the reader’s convenience we describe the steps in the proof.
First since the contact structure is stably parallelizable it follows that the tangent bundle of $M$ is stably trivial (forgetting the complex structure) so that the classifying map of its tangent bundle factors through a point:
@@
By surgery as in [Kre99, Lemma 2] we can add handles to $M\times I$ to inductively kill homotopy groups up to dimension $n$; let $W$ be the corresponding cobordism. Notice that the positive boundary of $W$ is a homotopy sphere by construction.
@@
Thus, since $TW$ is stably trivial as a real bundle, we can endow a stabilization of it with an almost complex structure, i.e. we found a stable almost complex structure on $W$. The conclusion of the lemma then directly follows from the fact that, according to [BCS14, Lemma 2.16], one can destabilize the stable almost complex structure on $W$ that we just found to a genuine almost complex structure which agrees (up to homotopy) with the given one on the negative end. ∎
As we already seen in Lemma 5.4, it is important to keep track of the Conley-Zehnder indices for non-contractible orbits to understand the indices of the Reeb orbits that become contractible on the spheres obtained via the flexible surgeries. For this, assume the contact manifold $Y$ has vanishing rational first Chern class. If we fix a complex trivialisation $\tau$ of the contact structure (or sum of copies of the contact structure, if the first Chern class in non-zero) over the $2$-skeleton, or equivalently a trivialisation of the complex determinant bundle associated to the contact structure, then we can speak of the Conley-Zehnder indices for any orbits with respect to $\tau$, which we will denote by $CZ_{\tau}(\gamma)$. One can then define notions of index positivity and ADC, that take into account all periodic orbits and not just those that are contractible. In this case we will say that $(Y,\tau)$ is generalized ADC following [Zho22b, Definition 3.1].
@@
Suppose that $V$ is Liouville with $c_{1}(V)$ torsion, then the contact boundary $\partial(V\times\mathbb{D}^{2})$ is generalized ADC with respect to any trivialization $\tau$.
@@
This proposition was essentially established in [Zho21b, Theorem 6.3]; we give some further details here for the readers’ ease.
We first choose a trivialisation $\tau^{\prime}_{0}$ of $\det_{\mathbb{C}}\oplus^{N}TV$ for some $N\in\mathbb{N}_{+}$ over $V$, which induces a trivialisation $\tau_{0}$ of $\det_{\mathbb{C}}\oplus^{N}T(V\times\mathbb{D}^{2})$ over $V\times\mathbb{D}^{2}$. Following the proof of [Zho21b, Theorem 6.3] and the notations therein, $\partial(V\times\mathbb{D}^{2})$ can then be decomposed into $Y^{2}_{\rho,f,g}$, diffeomorphic to $V\times S^{1}$, and $Y^{1}_{\rho,f,g}$, diffeomorphic to $\partial V\times\mathbb{D}^{2}$. The Reeb orbits (with period up to some threshold) on $Y^{2}_{\rho,f,g}$ are contractible and have positive SFT degrees. The Reeb orbits on $Y^{1}_{\rho,f,g}$ are described in [Zho21b, Proposition 6.7], where the Conley-Zehnder indices w.r.t. $\tau_{0}$ are computed using the Conley-Zehnder indices of the $\partial V$-component w.r.t. $\tau^{\prime}_{0}$. In particular, those orbits have arbitrarily high Conley-Zehnder indices w.r.t. $\tau_{0}$.
Lastly, for any trivialisation $\tau$ of $\det_{\mathbb{C}}\oplus^{N}T(V\times\mathbb{D}^{2})|_{\partial(V\times\mathbb{D}^{2})}$, the difference from the Conley-Zehnder index of $\gamma$ w.r.t. $\tau$ and $\tau_{0}$ is given by $\langle\,\tau-\tau_{0},\gamma\,\rangle$, where we view $\tau-\tau_{0}$ as a class in $H^{1}(\partial(V\times\mathbb{D}^{2}))$. Then the proof of [Zho21b, Theorem 6.3] implies that $\partial(V\times\mathbb{D}^{2})$ is generalized ADC w.r.t. $\tau$. ∎
@@
We let $V$ be a Liouville domain as constructed by Massot-Niederkrüger-Wendl [MNW13], so that $V\cong N\times I$ and $N^{2n-1}=G/\Gamma$ is a quotient of a solvable Lie group by a co-compact lattice. More precisely, the contact structures on the boundary components of $V$ come from left invariant structures on $G$, and are hence parallelizable. The same holds for the (symplectic) tangent bundle on $V$, and thus for $V\times\mathbb{D}^{2}$.
@@
Now set $M=\partial(V\times\mathbb{D}^{2})$ to be the contact boundary of the product. The underlying almost contact structure is stably trivial and we can apply Lemma 6.1 to obtain an almost Weinstein cobordism $W$ to some homotopy sphere. Combining Proposition 6.2 with [Laz20a], and realising $W$ as a flexible Weinstein cobordism using [CE12], we obtain an ADC contact structure on the convex end, which is a homotopy sphere $(\Sigma,\xi_{\Sigma})$. Moreover, this has a natural Liouville filling obtained by attaching Weinstein handles to $V\times\mathbb{D}^{2}$, which in particular has trivial symplectic cohomology by [Cie02b, BEE12]. The cohomology of the resulting filling is non-trivial in degree $12n-$ 1 corresponding to the fundamental class of $N$.
@@
Using the ADC property and the vanishing of its symplectic cohomology, according to [Zho21b, Corollary B] the cohomology of any filling of $(\Sigma,\xi_{\Sigma})$ with vanishing first Chern class, and in particular any Weinstein filling, has cohomology identical to the cohomology of the natural Liouville filling, and hence has non-trivial cohomology in degree $12n-$ 1, which is bigger than $n+1$ under our assumption of $n\geq 3$. So we deduce that $(\Sigma,\xi_{\Sigma})$ is Liouville, but not Weinstein, fillable.
@@
Now since the set of smooth, oriented homotopy spheres with the operation of connect sum forms a finite group, up to taking self connected sums we can assume that $\Sigma\cong\mathbb{S}^{2n+1}$ is diffeomorphic to the standard sphere, which then also has an ADC contact structure with a Liouville filling with vanishing symplectic cohomology that is homologically not Weinstein. In other words, we have found a contact structure on the standard sphere which is Liouville but not Weinstein fillable.
@@
To get the standard almost contact structure we use the fact that all stably trivial almost contact structures on a standard sphere can be realised via contact structures (cf. [BCS14, Lemma 2.17]) which are Weinstein fillable. Taking connected sum of these with the previously constructed Liouville non-Weinstein fillable structures then gives one structure as in the statement of Theorem C. ∎
@@
One can now easily get infinitely many distinct contact structures as in Theorem C by taking the connected sum with the infinitely many examples of homotopically standard and flexibly fillable contact structures on spheres from [Laz20a]. Arguing mutatis mutandis we also obtain Theorem E from the Introduction. We also point out that the $5$-dimensional case remains open since, for all known examples of four-dimensional Liouville domains $V$ with disconnected boundary, the manifold $V\times\mathbb{D}^{2}$ is Weinstein by Breen-Christian [BC21].

### 14 Mar 2026 v5

@@ 5.2. Abundance of exotic structures on spheres
When $n\geq 5$, there exists a Weinstein fillable contact structure $(\mathbb{S}^{2n+1},\xi^{\prime})$ homotopic to the standard almost contact structure, which is $1$-ADC, and such that $0<\dim SH^{2n-3}_{+}(\mathbb{S}^{2n+1},\xi^{\prime})<\infty$.
The construction of $\xi^{\prime}$ in Proposition 5.5 is based on the description given in [KvK16, §5.3.1] of the Reeb orbits of an explicit contact form on the Brieskorn manifold $\Sigma(a_{0},\ldots,a_{n}):=\{z_{0}^{a_{0}}+\ldots+z_{n}^{a_{n}}=1\}\cap S^{2n+1}$. We first set up some terminology that we will need later on in the description of the Reeb flow from [KvK16, §5.3.1].
A positive real number $T$ is called a principal period if there exists $I_{T}\subset I_{n}=\{0,\ldots,n\}$ such that:
  1. (1)
    $a_{i}|T$ if and only if $i\in I_{T}$;
  2. (2)
    $T=\operatorname{lcm}\{a_{i}\mid i\in I_{T}\}$;
  3. (3)
    $|I_{T}|\geq 2$.
Then, according to [KvK16, §5.3.1], the periods of the Reeb flow are of the form $NT$, for some $N\in\mathbb{N}_{+}$ and a principal period $T$. For a period $A$, there is a principal period $T$ such that $A=NT$ and there is no other principal period $T^{\prime}$ that divides $NT$ and is divided by $T$. There is then a family of dimension $2|I_{T}|-3$ (diffeomorphic to the Brieskorn manifold with coefficients given by the sequence $I_{T}$) of parametrized Reeb orbits with period $NT$, which, after taking the quotient by the action given by $S^{1}$-reparameterization becomes $2|I_{T}|-4$ dimensional. The generalized Conley-Zehnder index of such a family is given by
$2\sum_{i\in I_{T}}\frac{NT}{a_{i}}+2\sum_{i\notin I_{T}}\left\lfloor\frac{NT}{a_{i}}\right\rfloor+|I_{n}|-|I_{T}|-2NT.$
If we perturb the contact form into a non-degenerate one using a Morse function on the critical manifold following [Bou02b, Lemma 2.3], the minimal Conley-Zehnder index of the non-degenerate orbits coming from this family is given by the generalized Conley-Zehnder index minus half of the dimension of the family itself, i.e.
$2\sum_{i\in I_{T}}\frac{NT}{a_{i}}+2\sum_{i\notin I_{T}}\left\lfloor\frac{NT}{a_{i}}\right\rfloor+|I_{n}|-2|I_{T}|-2NT+2.$ (11)
Now, for each $A$, there might be several such principal periods $T$. Each will give rise to a Morse-Bott non-degenerate family of Reeb orbits, and they describe all of the Reeb orbits.
Proof of Proposition 5.5.
@@
We consider the Brieskorn sphere $Y:=\Sigma(n,\ldots,n,n+1,p)$, where $p$ is a prime number with $p\gg n$. The minimal Conley-Zehnder index of a small perturbation of the standard contact form on $Y$ is 4n4-n, obtained from the expression in (11) above as $|I_{n+1}|=n+2$, $|I_{T}|=n$, $N=1$, $T=n$ in our setup. The SFT degrees are then at least $2$. The minimal Conley-Zehnder indices of other Morse-Bott families of Reeb orbits of the natural contact form (after suitable perturbation) are 6n6-n (given by $N=2$, $T=n$ in (11)) as $p\gg n$, and so they have SFT degree $4$. As a consequence, we have that $Y$ is $1$-ADC. Moreover, by the Morse-Bott spectral sequence (see e.g. [KvK16, Theorem 5.4.]) [^8] Note that [KvK16] uses homological grading by $\mu_{CZ}$ while we use cohomological grading $n+1-\mu_{CZ}$ and the index gap above, we deduce that
$\dim SH_{+}^{(n+1)-(4-n)}(Y)=\dim SH_{+}^{2n-3}(Y)=1\,$,
which is generated by the family of minimal Conley-Zehnder index. Now, by [KvK16, Proposition 3.6], $Y$ is homeomorphic to a sphere. Then there exists $K$, such that the $K$-th iterated self-connected sum $^{K}Y$ is diffeomorphic to the standard sphere, as the group of smooth structures on $S^{k}$ is finite for $k\neq 4$, see [KM63]. Since $\dim Y\geq 11$, the contact sum preserves the $1$-ADC property: indeed, by the proof of [Laz20a, Theorem 3.14, 3.15], an index $k$ subcritical handle attachment to a 2n12n-1 dimensional contact manifold preserves the $p$-ADC property for any $p\leq 2n-3-k$. Moreover, by [CO18, Theorem 7.1 and Proposition 9.19] we have
$\dim SH_{+}^{2n-3}(\#^{K}_{i=1}Y)=\dim(\bigoplus^{K}_{i=1}SH^{2n-3}_{+}(Y))=K\,.$
To obtain the standard almost contact structure, we choose $Y^{\prime}$ to be the flexibly fillable sphere with the opposite homotopy class  [^9] To be precise, here “opposite” is to be interpreted as the inverse w.r.t. the natural group structure on the space of almost contact structures on the sphere given by the connected sum operation. of almost contact structures as that of $^{K}Y$, as in [Zho22, Corollary 5.19]. Then $Y^{\prime}\#^{K}Y$ is a standard sphere with the standard almost contact structure, and it is $1$-ADC. Moreover, $SH^{2n-3}_{+}(Y^{\prime})=0$ when $n\geq 5$. Hence, we may take $Y^{\prime}\#^{K}Y$ as the desired contact sphere. ∎
@@
One first realizes the given almost contact structure as the contact boundary $(M,\xi_{\textit{flex}})$ of a flexible Weinstein domain using Eliashberg’s $h$-principle [CE12]. For $i\in\mathbb{N}$, we then consider the iterated connected sums
$Y_{i}:=(\mathbb{S}^{2n+1},\xi_{ex})\#(M,\xi_{\textit{flex}})\#^{i}(\mathbb{S}^{2n+1},\xi^{\prime})$
with the contact sphere $(\mathbb{S}^{2n+1},\xi_{ex})$ from Theorem 5.1 and $(\mathbb{S}^{2n+1},\xi^{\prime})$ from Proposition 5.5. Notice that $M$ is $1$-ADC, as it is flexibly fillable [Laz20a], and that $SH^{*}_{+}(M)$ is supported in degree smaller than $n+1$, in view of the vanishing of symplectic homology for flexible Weinstein domains. Since $SH_{+}^{*}(\mathbb{S}^{2n+1},\xi_{ex})$ is supported in degrees at most $n+1$ by Lemma 5.4, then if 2n3>n+12n-3>n+1, i.e. $n\geq 5$, we have $\dim SH^{2n-3}_{+}(Y_{i})=N\cdot i$, where $N=\dim SH^{2n-3}_{+}(\mathbb{S}^{2n+1},\xi^{\prime})$, by [CO18, Theorem 7.1, Proposition 9.19]. As a consequence, the $Y_{i}$’s are pairwise distinguished by $SH^{*}_{+}$. Moreover, they are tight and not strongly fillable by the same arguments as in the proof of Theorem A. ∎
@@
When $n=3$, $(\mathbb{S}^{2n+1},\xi_{ex})$ is not $1$-ADC, which costs us the well-definedness of the positive symplectic cohomology. When $n=4$, it is not easy to find a degree such that positive symplectic cohomology is different for varying summands in the connected sum. In those cases, we can use positive symplectic cohomology for augmentations in [Zho23a, §3.2.3] [^10] The positive symplectic cohomology of $1$-ADC manifolds is the special case for the trivial augmentation., which was motivated by the work of Bourgeois-Oancea [BO09a]. Then we can obtain a “contact invariant” by enumerating through all possible augmentations. However, the current status of the contact homology [Par19, BH23] is not enough to guarantee that such a contact invariant is independent of various choices. This being said, assuming the invariance of such a theory, one can upgrade Theorem D to cover the cases of $n=3,4$.
@@ 6. Liouville fillable contact spheres without Weinstein fillings
We now explain how to use surgery techniques to deduce the existence of Liouville fillable contact structures that are not Weinstein fillable on spheres. This requires a refinement of the surgery arguments used above, and we begin with the following:
@@
Let $(M^{2n+1},\xi)$ be a contact manifold of dimension at least $5$ whose underlying almost contact structure is stably trivial (as an almost contact structure). Then there is an almost Weinstein cobordism $W$ from $(M,\xi)$ to a homotopy sphere $\Sigma$ (with some almost contact structure).
@@
This follows immediately from results in [BCS14], building on the work of Kreck [Kre99]. For the reader’s convenience, we describe the steps in the proof.
First, since the contact structure is stably parallelizable, it follows that the tangent bundle of $M$ is stably trivial (forgetting the complex structure) so that the classifying map of its tangent bundle factors through a point:
@@
By surgery as in [Kre99, Lemma 2], we can add handles to $M\times I$ to inductively kill homotopy groups up to dimension $n$; let $W$ be the corresponding cobordism. Notice that the positive boundary of $W$ is a homotopy sphere by construction.
@@
Thus, since $TW$ is stably trivial as a real bundle, we can endow a stabilization of it with an almost complex structure, i.e. we found a stable almost complex structure on $W$. The conclusion of the lemma then directly follows from the fact that, according to [BCS14, Lemma 2.16], one can destabilize the stable almost complex structure on $W$ that we just found to a genuine almost complex structure which agrees (up to homotopy) with the given one on the negative end. Since this cobordism was obtained by attaching handles of index at most $n$ it is, in particular, almost Weinstein. ∎
As we have already seen in Lemma 5.4, it is important to keep track of the Conley-Zehnder indices for non-contractible orbits to understand the indices of the Reeb orbits that become contractible on the spheres obtained via the flexible surgeries. For this, assume the contact manifold $Y$ has vanishing rational first Chern class. If we fix a complex trivialisation $\tau$ of the contact structure (or of the sum of the right number of copies of the contact structure, if the first Chern class is a non-zero torsion) over the $2$-skeleton, or equivalently a trivialisation of the complex determinant bundle associated to the contact structure, then we can speak of the Conley-Zehnder indices for any orbits with respect to $\tau$, which we will denote by $\mu_{CZ}^{\tau}(\gamma)$. One can then define notions of index positivity and ADC that take into account all periodic orbits and not just those that are contractible. This leads to the following concept of $\pi_{1}$-sensitive ADC manifolds introduced in [Zho24, Definition 3.5].
Definition 6.2.
Let $(Y,\xi)$ be a contact manifold such that $c_{1}^{\mathbb{Q}}(\xi)=0$. Let $\tau$ be a trivialization of $\det_{\mathbb{C}}\left(\oplus^{N}\xi\right)$ for some $N\in\mathbb{N}_{+}$. We say $(Y,\xi,\tau)$ is $\pi_{1}$-sensitive ADC if there exist contact forms $\alpha_{1}>\alpha_{2}>\ldots$, and positive real numbers $D_{1}<D_{2}<\ldots$ converging to infinity, such that all Reeb orbits of $\alpha_{i}$ of period up to $D_{i}$ are non-degenerate and have rational SFT grading $\mu_{CZ}^{\tau}(\gamma)+n-3>0$.
@@
Suppose that $V$ is Liouville with $c_{1}(V)$ torsion, then the contact boundary $\partial(V\times\mathbb{D}^{2})$ is $\pi_{1}$-sensitive ADC with respect to any trivialization $\tau$.
@@
This proposition was essentially established in [Zho21b, Theorem 6.3]; we give some further details here for completeness.
We first choose a trivialisation $\tau^{\prime}_{0}$ of $\det_{\mathbb{C}}\oplus^{N}TV$ for some $N\in\mathbb{N}_{+}$ over $V$, which induces a trivialisation $\tau_{0}$ of $\det_{\mathbb{C}}\oplus^{N}T(V\times\mathbb{D}^{2})$ over $V\times\mathbb{D}^{2}$. Following the proof of [Zho21b, Theorem 6.3] and the notations therein, $\partial(V\times\mathbb{D}^{2})$ can then be decomposed into $Y^{2}_{\rho,f,g}$, diffeomorphic to $V\times S^{1}$, and $Y^{1}_{\rho,f,g}$, diffeomorphic to $\partial V\times\mathbb{D}^{2}$. We obtain a sequence of decreasing contact forms by choosing $\rho,f,g$. Here $f,g$ are perturbation functions, and $\rho$ is a positive number measuring the size of $V$ compared to $\mathbb{D}^{2}$, i.e. a large $\rho$ results in a contact form that is “thin in the $\mathbb{D}^{2}$ direction” (such as what happens for the contact form at the boundary of a thin ellipsoid), see [Zho21b, Proof of Theorem 6.3] for details. The Reeb orbits (with period up to some threshold) on $Y^{2}_{\rho,f,g}$ are contractible and have positive SFT degrees. Since those orbits are contractible, the SFT degree is independent of the trivialization [GZ21, Proposition 3.8]. The Reeb orbits on $Y^{1}_{\rho,f,g}$ are described in [Zho21b, Proposition 6.7], where the Conley-Zehnder indices w.r.t. $\tau_{0}$ are computed using the Conley-Zehnder indices of the $\partial V$-component w.r.t. $\tau^{\prime}_{0}$. In particular, in this sequence of contact forms, those orbits have arbitrarily high Conley-Zehnder indices w.r.t. $\tau_{0}$. Heuristically speaking, those orbits are similar to the orbits rotating along the long axis of a very thin ellipsoid.
Lastly, for any trivialization $\tau$ of $\det_{\mathbb{C}}\oplus^{N}T(V\times\mathbb{D}^{2})|_{\partial(V\times\mathbb{D}^{2})}$, the difference from the Conley-Zehnder index of $\gamma$ w.r.t. $\tau$ and $\tau_{0}$ is given by $\frac{2}{N}\langle\,\tau-\tau_{0},\gamma\,\rangle$, where we view $\tau-\tau_{0}$ as a class in $H^{1}(\partial(V\times\mathbb{D}^{2}))$, see [GZ21, §3.3.2]. Therefore, the error is a bounded number for any fixed homology class of $\gamma$. On the other hand, for any fixed homology class, Reeb orbits on $Y^{1}_{\rho,f,g}$ in that class have arbitrarily high Conley-Zehnder indices w.r.t. $\tau_{0}$ from the proof of [Zho21b, Theorem 6.3], this implies that $\partial(V\times\mathbb{D}^{2})$ is $\pi_{1}$-sensitive ADC w.r.t. $\tau$. ∎
Lazarev [Laz20a, Theorems 3.14, 3.15] proved that the ADC property is preserved under subcritical and flexible surgeries as long as the vanishing of the rational first Chern class condition holds. The case of attaching a $2$-handle [Laz20a, Theorem 3.14] has extra assumptions since Lazarev only considered contractible Reeb orbits. In the case of $\pi_{1}$-sensitive ADC manifolds, [Laz20a, Theorems 3.14, 3.15] holds with identical proofs as long as the trivializations are compatible:
Proposition 6.4.
Let $W$ be a flexible cobordism from $Y_{-}$ to $Y_{+}$ with $c_{1}^{\mathbb{Q}}(TW)=0$. Let $\tau$ be a trivialization of $\det_{\mathbb{C}}\oplus^{N}TW$ whose restriction to $Y_{\pm}$ is denoted by $\tau_{\pm}$. If $Y_{-}$ is $\pi_{1}$-sensitive ADC w.r.t. $\tau_{-}$, $Y_{+}$ is $\pi_{1}$-sensitive ADC w.r.t. $\tau_{+}$.
@@
We let $V$ be a Liouville domain as constructed by Massot, Niederkrüger, and Wendl [MNW13], so that $V\cong N\times I$ and $N^{2n-1}=G/\Gamma$ is a quotient of a solvable Lie group by a co-compact lattice. More precisely, the contact structures on the boundary components of $V$ come from left-invariant structures on $G$, and are hence parallelizable. The same holds for the (symplectic) tangent bundle on $V$, and thus for $V\times\mathbb{D}^{2}$.
@@
Now set $M=\partial(V\times\mathbb{D}^{2})$ to be the contact boundary of the product. The underlying almost contact structure is stably trivial, and we can apply Lemma 6.1 to obtain an almost Weinstein cobordism $W$ to some homotopy sphere. Since the almost complex structure on $W$ is obtained from destabilizing a trivial complex structure, the first Chern class of $W$ is zero. Combining Proposition 6.3 with Proposition 6.4, and realizing $W$ as a flexible Weinstein cobordism using [CE12, Theorem 13.1], we obtain an ADC contact structure on the convex end, which is a homotopy sphere $(\Sigma,\xi_{\Sigma})$. Moreover, this has a natural Liouville filling obtained by attaching Weinstein handles to $V\times\mathbb{D}^{2}$, which in particular has trivial symplectic cohomology by [Cie02a, Theorem 1.11] and [BEE12, Theorem 5.6]. The cohomology of the resulting filling has a non-trivial element in degree 2n12n-1, which evaluates non-trivially on the fundamental class of $N$.
@@
Using the ADC property and the vanishing of its symplectic cohomology, according to [Zho21b, Corollary B] the cohomology of any Weinstein filling of $(\Sigma,\xi_{\Sigma})$ has cohomology identical to the cohomology of the natural Liouville filling, and hence has non-trivial cohomology in degree 2n12n-1, which is bigger than $n+1$ under our assumption of $n\geq 3$. So we deduce that $(\Sigma,\xi_{\Sigma})$ is Liouville, but not Weinstein, fillable. Such phenomena were observed in [Zho21b].
@@
Now since the set of smooth, oriented homotopy spheres with the operation of connected sum forms a finite group, up to taking self-connected sums we can assume that $\Sigma\cong\mathbb{S}^{2n+1}$ is diffeomorphic to the standard sphere, which then also has an ADC contact structure with a Liouville filling with vanishing symplectic cohomology that is homologically not Weinstein. In other words, we have found a contact structure on the standard sphere which is Liouville but not Weinstein fillable.
@@
To get the standard almost contact structure, we use the fact that all stably trivial almost contact structures on a standard sphere can be realised via contact structures (cf. [BCS14, Lemma 2.17]) which are flexibly fillable. Taking a connected sum of one of such contact structures with the previously constructed Liouville non-Weinstein fillable structures then gives one structure as in the statement of Theorem C. ∎
@@
One can now easily get infinitely many distinct contact structures as in Theorem C by taking the connected sum with the infinitely many examples of homotopically standard and flexibly fillable contact structures on spheres from [Laz20a]. Arguing mutatis mutandis, we also obtain Theorem E from the Introduction. We also point out that the $5$-dimensional case remains open since, for all known examples of four-dimensional Liouville domains $V$ with disconnected boundary, the manifold $V\times\mathbb{D}^{2}$ is Weinstein by Breen-Christian [BC25].